Suppose that your demand schedule for compact discs is as follows: PRICE Quantity Demanded $ 8 10 12 14 16 Use the midpoint method to calculate your price elasticity of demand as the price of compact discs 40 32 24 16 increases from $8 to $10.
Answer #1

P1 =$8

P2 =$10

Q1= 40

Q2 =32

By Midpoint method

Price Elasticity of demand = (% change in quantity demanded/ % change in price)

% change in quantity demanded = (32-40)/[(32+40)/2] 100 = (-8 / 36) 100 = (-0.22)(100) =22.22

% change in price = (10-8)/ [(10+8)/2] 100 =( 2/9)100 =(0.22)(100) =22.22

Elasticity of demand = (-22.22/22.22)= -1

The elasticity of demand is unit elastic as the price of CD's increases from $8 to $10.